OpenStudy (kearsten):

a student wrote the opposite of the expression(-4+x) as(x-4). what was his mistake? what is the correct answer?

zepdrix (zepdrix):

If by `opposite` you mean `additive inverse`, then the opposite of (-4+x) would be -(4+x) Since addition is commutative, a+b = b+a which is what is being illustrated in your example.

zepdrix (zepdrix):

-4 + x = x + -4

zepdrix (zepdrix):

Unless I'm misunderstanding what you meant by `opposite` +_+ Bahhh I dunno

OpenStudy (kearsten):

thnx